OpenStudy (anonymous):

A clothing designer enters a competition to show ten original outfits at a fashion show. If the designer has 12 outfits to choose from how many ways to arrange the order in which he will show ten of them.

OpenStudy (kropot72):

This requires calculation of permutations: \[12P10=\frac{12!}{(12-10)!}=\frac{12\times11\times10\times9\times8\times7\times6\times5\times4\times3}{2}=you\ can\ calculate\]

OpenStudy (anonymous):

I did everything but divide by 2. I didn't know that I had to do that but my answer was so large. I can't see your answer

OpenStudy (kropot72):

The formula for finding the number of permutations of n different things taken r at a time is: \[nPr=\frac{n!}{(n-r)!}\] Can you use a calculator to work out the answer and post it?

OpenStudy (anonymous):

239,500,800/2=119750400 is this right?

OpenStudy (kropot72):

\[\frac{12\times11\times10\times9\times8\times7\times6\times5\times4\times3\times2\times1}{2}=\frac{479001600}{2\times1}=239500800\] Sorry, the numerator of my first post left out (*2*1).